OpenStudy (anonymous):

Geometric mean help :S Find the three geometric means between 1/ 2 and 8.

OpenStudy (anonymous):

can you show me the steps as well?

OpenStudy (anonymous):

As far as I know Geometric mean of A and B: \[\sqrt{A*B}\]

OpenStudy (anonymous):

(1/2)(8) => b = 2

OpenStudy (anonymous):

im confused on the others..

OpenStudy (anonymous):

So 1/2 and 8 are suppose to have 3 geometric means?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't think that's possible.. 2 real numbers always have 1 geometric mean.

OpenStudy (anonymous):

what about 1,2,4?

OpenStudy (anonymous):

\[\sqrt[3]{1 * 2 * 4} = \sqrt[3]{8} = 2\]

OpenStudy (anonymous):

The geometric mean is 1,2,4 because (1/2)(8) = 2 ?

OpenStudy (anonymous):

(1/2)(8) is 4 not 2. radical 4 is 2. So geometric mean of 1/2 and 8 is 2.

OpenStudy (anonymous):

okay so its impossible for these numbers to have three geometric means.

OpenStudy (anonymous):

yes